6. What mass of water (in kg) at 1000°C could be energy: 5.00E kg vaporized with 2.70E cal of 7. How many joules (J) of energy are released when 6.80x103g of steam at 100.0 C are completely frozen to ice at 0.0°C 2.05EJ 8. Convert the H value for water (80 cal/g) to units of Jlimal: (HINT: You will need to use the molar mass of water) 6025 J/mol [9] How much energy (in J) is required to completely melt 205.0 mol of ice at 0.o C: 1.235e s (sol.) Compound! ) M.P s (liq-) B.P s (gas) 334 61.4 H20 2.09 0.560 62 -39 961 4.184 1.070 0.138 100 760 357 2260 2025 294 2355 .97 0.671 0.104 Hg 0.217 105 at 100℃. The specific heat (s) of steam ncreases slighty with roeasng temperature For example, te Cp of steam at 200C is 2.00. [10] How much heat is needed to raise the C totemperature125°C? of a [11] How much heat is needed to raise the temperature of 85 g of potassium from 25°C to 2,500°C?
Answer #1

(7)

Heat released in this complete conversion = 6800*2260 + 6800*4.184*(100-0) + 6800*334 = 20484320 J = 2.05*107 J

(8)

1 cal = 4.184 J

1 mol water = 18 g

So,

80 cal/g = 80*(4.184*18) = 6024.96 J/mol = 6025 J/mol

Hope this helps !